OpenStudy (anonymous):

-(m+2)=3+m step by step plz

OpenStudy (anonymous):

distributue the negative into the parentehese: -m-2=3+m Then we can add m to both sides: -2=3+2m Next subtract 3 from both sides:-5=2m Finally divide both sides by 2 and get:-5/2=m

OpenStudy (anonymous):

that is confussing

OpenStudy (anonymous):

-(m+2)=3+m Distribute the negtaive sign into the parentheses:-m-2=3+m Subtract m from both sides: -m-2-m=3+m-m Simplify:-2m-2=3 Add 2 to both sides:-2m-2+2=3+2 Simplify:-2m=5 Now divide both sides by -2:m=-5/2